OpenStudy (anonymous):

One term of an arithmetic sequence is 73, and the next term is 90. The fourteenth term of the sequence is 209. Please help find the first term of the sequence, and explain how you found it??

OpenStudy (anonymous):

The difference between successive terms is 90 - 73 = 17 Fourteen terms would occupy 14 × 17 = 238 So the first term is: 209 - 238 = -29